### v1

# Computational and Nonlinear Dynamics

[…]
Computational and Nonlinear Dynamics
[…]
Auburn University,
Auburn, AL 36849,
email : mzm0390@auburn.edu
[…]
Dan B. Marghitu Department of Mechanical Engineering,
Auburn University,
Auburn, AL 36849,
email : marghdb@auburn.edu
Asymptotic Behavior of an Unforced Duhem-Type Hysteretic Oscillator
(5 janvier 2026)
Résumé
The article describes fundamental analytical properties of an unforced mechanical oscillator with a Duhem-type viscoelastoplastic hysteretic element. These properties include global existence of solutions, uniqueness of solutions, and convergence of each solution to an equilibrium point.
keywords:
Multibody System Dynamics, Nonlinear Dynamical Systems
1 Introduction
The Duhem models constitute a class of differential models of hysteresis that are suitable for the description of a variety of physical phenomena, including dry friction, elastoplastic materials, and magnetization. The models are usually attributed to Pierre Duhem [Duhem (1896-1902), as cited in ikhouane_survey_2018]. However, the interest in the models has surged only in the second half of the twentieth century [everett_general_1954, chua_mathematical_1971, chua_generalized_1972, pokrovskii_theory_1973, visintin_continuity_1983, coleman_constitutive_1986, coleman_class_1987, krasnoselskii_systems_1989, visintin_differential_1994, oh_modeling_2003, oh_analysis_2003, oh_semilinear_2005, padthe_counterclockwise_2005, jayawardhana_sufficient_2009, jayawardhana_dissipativity_2011, jayawardhana_stability_2012, ouyang_stability_2012, naser_consistency_2013, naser_characterization_2013, vasquez-beltran_modeling_2023]. Specializations of the models include the Dahl friction model [dahl_solid_1968], the LuGre friction model [canudas_de_wit_new_1995], and the Bouc-Wen model [bouc_forced_1967, bouc_modemathematique_1971, wen_method_1976]. Refs. [visintin_differential_1994] and [ikhouane_survey_2018, ikhouane_erratum_2018] provide further general information about the Duhem models. Refs. [krasnoselskii_systems_1989, mayergoyz_mathematical_1991, visintin_differential_1994, mielke_rate-independent_2015] provide further information about differential models of hysteresis in general.
The following form of a Duhem model is essentially equivalent [^1] Apart from syntactic discrepancies, Ref. [oh_analysis_2003] employs a different set of regularity conditions for $f$, $g$, and $h$. to the form employed in Ref. [oh_analysis_2003] : [^2] See Appendix A for notation and conventions.
$\begin{cases}\dot{x}=u\\ \dot{z}=f(x,z)g(u)\\ y=-h(x,z)\\ \begin{matrix}x(0)=x_{0}&z(0)=z_{0}\end{matrix}\end{cases}$ (1)
where $x\in\mathbb{R}$ and $z\in\mathbb{R}^{n}$ are state variables, $u\in\mathbb{R}$ is an input variable, $y\in\mathbb{R}$ is an output variable, $x_{0},z_{0}\in\mathbb{R}$ are parameters, $f:\mathbb{R}\times\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n\times p}$, $g:\mathbb{R}\longrightarrow\mathbb{R}^{p}$, $h:\mathbb{R}\times\mathbb{R}^{n}\longrightarrow\mathbb{R}$ are continuous functions, and $n,p\in\mathbb{Z}_{\geq 1}$.
In this article, the following specialization of the model given by Eq. (1) will be considered :
$\begin{cases}\dot{x}=u\\ \dot{z}=u+f_{1}(z)g_{1}(u)+f_{2}(z)g_{2}(u)\\ y=-h_{1}(x)-h_{2}(z)\\ \begin{matrix}x(0)=x_{0}&z(0)=z_{0}\end{matrix}\end{cases}$ (2)
Here, $x,z\in\mathbb{R}$ are state variables, $u\in\mathbb{R}$ is an input variable, $y\in\mathbb{R}$ is an output variable, $f_{1},f_{2},g_{1},g_{2}:\mathbb{R}\longrightarrow\mathbb{R}$ are locally Lipschitz continuous functions, and $h_{1},h_{2}:\mathbb{R}\longrightarrow\mathbb{R}$ are strictly increasing locally Lipschitz continuous functions (and, therefore, homeomorphisms) such that
$f_{1}(0)=f_{2}(0)=g_{1}(0)=g_{2}(0)=h_{1}(0)=h_{2}(0)=0$ (3)
$\forall z\in\mathbb{R}_{<0}.\;0<f_{1}(z)\wedge f_{2}(z)<0$ (4)
$\forall z\in\mathbb{R}_{>0}.\;f_{1}(z)<0\wedge 0<f_{2}(z)$ (5)
$\forall v\in\mathbb{R}_{>0}.\;0<g_{1}(v)\wedge g_{2}(v)=0$ (6)
$\forall v\in\mathbb{R}_{<0}.\;g_{1}(v)=0\wedge g_{2}(v)<0$ (7)
The following model represents an unforced mechanical oscillator (e.g., see Ref. [mickens_truly_2010]) with a viscoelastoplastic Duhem-type element and (optionally) additional viscous damping : [^3] For simplicity, the parameter that represents the inertia of the system is ignored. The inertial properties of the system can be taken into account by rescaling or making appropriate adjustments to $h_{1}$, $h_{2}$, and $c$.
$\begin{cases}\dot{x}=v\\ \dot{z}=v+f_{1}(z)g_{1}(v)+f_{2}(z)g_{2}(v)\\ \dot{v}=-h_{1}(x)-h_{2}(z)-c(x,z,v)\\ \begin{matrix}x(0)=x_{0}&z(0)=z_{0}&v(0)=v_{0}\end{matrix}\end{cases}$ (8)
Here, $x,z,v\in\mathbb{R}$ are state variables, and $x_{0},z_{0},v_{0}\in\mathbb{R}$ are parameters, $c:\mathbb{R}^{3}\longrightarrow\mathbb{R}$ is a locally Lipschitz continuous function such that
$\forall x,z\in\mathbb{R}.\;c(x,z,0)=0$ (9)
$\forall x,z,v\in\mathbb{R}.\;0\leq vc(x,z,v)$ (10)
The article addresses the problem of the asymptotic behavior of solutions of the system given by Eq. (8). More specifically, it is shown that every trajectory of the system converges to an equilibrium point. This result is a generalization of a result about the asymptotic behavior of the Bouc-Wen model of class I that was previously presented in Refs. [ikhouane_dynamic_2007, ikhouane_systems_2007]. It should be noted that other closely related problems were considered in [anderssen_global_1998, jayawardhana_stability_2012, ouyang_stability_2012, ouyang_absolute_2014].
The remainder of the article is organized as follows :
  • Section 2 provides proofs of global existence, uniqueness and boundedness of solutions of Eq. (8).
  • Section 3 provides a proof of convergence of each solution of Eq. (8) to an equilibrium point.
  • Section 4 presents an application of the results developed in Section 3 to a Bouc-Wen oscillator.
  • Section 5 provides conclusions and recommendations.
  • Appendix A describes the mathematical conventions and provides the proofs of secondary results.
2 Global Existence, Uniqueness, Boundedness
Define
$\mathcal{E}\triangleq\left\{(x,z,v)\in\mathbb{R}^{3}:h_{1}(x)=-h_{2}(z)\wedge v=0\right\}$ (11)
Then,
Proposition 2.1.
$\mathcal{E}$ is the set of all equilibrium points of the system described by Eq. (8).
Démonstration.
Suppose that $(x_{e},z_{e},v_{e})\in\mathcal{E}$. Note that $v_{e}=0$. Then, since $g_{1}(v_{e})=g_{2}(v_{e})=0$, $\dot{z}=0$ and $\dot{x}=0$. Furthermore, $c(x_{e},z_{e},v_{e})=0$ and $h_{1}(x_{e})-h_{2}(z_{e})=0$. Therefore, $\dot{v}=0$. Thus, $(x_{e},z_{e},v_{e})$ is an equilibrium point of the system described by Eq. (8). Suppose that $(x_{e},z_{e},v_{e})$ is an equilibrium point of the system described by Eq. (8). Then, since $\dot{x}=0$, $v_{e}=0$. Thus, $c(x_{e},z_{e},v_{e})=0$. Therefore, since $\dot{v}=0$, $h_{1}(x_{e})-h_{2}(z_{e})=0$. Hence, $(x_{e},z_{e},v_{e})\in\mathcal{E}$. ∎
Consider the following Lyapunov function candidate :
$\mathcal{V}(x,z,v)=\int_{0}^{x}h_{1}(s)ds+\int_{0}^{z}h_{2}(s)ds+\frac{1}{2}v^{2}$ (12)
The following lemmas show that $\mathcal{V}$ is positive definite and radially unbounded :
Lemma 2.2.
$0\leq\mathcal{V}(x,z,v)$ for all $x,z,v\in\mathbb{R}$ ; $\mathcal{V}(x,z,v)=0$ if and only if $x=z=v=0$.
Démonstration.
Note that $\leq\int_{0}^{u}h_{1}(s)ds, 00$$\leq\int_{0}^{u}h_{2}(s)ds$, and $0\leq(1/2)u^{2}$ for all $u\in\mathbb{R}$. Thus, $0\leq\mathcal{V}(x,z,v)$. It can be verified by substitution that $\mathcal{V}(0)=0$. Suppose that $\mathcal{V}(x,z,v)=0$. Note that $\int_{0}^{u}h_{1}(s)ds>0$ and $\int_{0}^{u}h_{2}(s)ds>0$ for all $u\in\mathbb{R}\setminus\{0\}$. Thus, $\int_{0}^{x}h_{1}(s)ds=0$, $\int_{0}^{z}h_{2}(s)ds=0$ and $\frac{1}{2}v^{2}=0$. Therefore, $x=z=v=0$. ∎
Lemma 2.3.
$\mathcal{V}$ is radially unbounded, that is, $\mathcal{V}(x,z,v)\rightarrow+\infty$ as $\lVert(x,z,v)\rVert_{\infty}\rightarrow+\infty$.
Démonstration.
[…]
It can be verified that
$\dot{\mathcal{V}}(x,z,v)=h_{2}(z)f_{1}(z)g_{1}(v)+h_{2}(z)f_{2}(z)g_{2}(v)-vc(x,z,v)$ (13)
Lemma 2.4.
$\dot{\mathcal{V}}(x,z,v)\leq 0$ for all $x,z,v\in\mathbb{R}$.
Démonstration.
Fix $x,z,v\in\mathbb{R}$. Note that $vc(x,z,v)\leq 0$. Note also that $h_{2}(z)f_{1}(z)\leq 0$ and $0\leq h_{2}(z)f_{2}(z)$ while $0\leq g_{1}(v)$ and $g_{2}(v)\leq 0$. Thus, $h_{2}(z)f_{1}(z)g_{1}(v)\leq 0$ and $h_{2}(z)f_{2}(z)g_{2}(v)\leq 0$. Thus, $\dot{\mathcal{V}}(x,z,v)\leq 0$ as a sum of nonpositive terms. ∎
Proposition 2.5.
The solutions of the system given by Eq. (8) are equibounded.
Démonstration.
Noting that $\mathcal{V}$ is continuously differentiable, the proof follows from Lemmas 2.2, 2.3, and 2.4 by, for example, Theorem 8.7 in Ref. [yoshizawa_stability_1975]. ∎
Remark.
See, e.g., Ref. [kellett_compendium_2014] for a description of a relationship between the radially unbounded positive definite functions and class $\mathcal{K}_{\infty}$ functions.
Theorem 2.6.
There exists a unique solution of the IVP given by Eq. (8) for every initial condition with $x_{0},z_{0},v_{0}\in\mathbb{R}$ on any time interval $[0,T)$ with $T\in\mathbb{R}_{>0}\cup\left\{+\infty\right\}$.
Démonstration.
Taking into account that the state function of the system described by Eq. (8) is locally Lipschitz continuous, the solutions are unique and exist on a non-empty maximal interval of existence (e.g., see Theorem 54 in Ref. [sontag_mathematical_1998]). Therefore, taking into account Proposition 2.5, by the theorem on the extendability of the solutions (e.g., see Proposition C.3.6 in Ref. [sontag_mathematical_1998]), each solution can be extended to a unique solution on $0,+\infty)$. ∎
3 Stability and Convergence
Proposition 3.1.
The origin $0\in\mathbb{R}^{3}$ is a Lyapunov stable equilibrium point of the system given by Eq. (8).
Démonstration.
That the origin is an equilibrium point can be verified by substitution into Eq. (8). Noting that $\mathcal{V}$ is continuously differentiable, Lyapunov stability of the origin follows from Lemmas 2.2, 2.3, and 2.4 by, for example, Theorem 4.1 in Ref. [khalil_nonlinear_2002]. ∎
Define
$\mathcal{M}_{z}\triangleq\left\{(x,z,v)\in\mathbb{R}^{3}:z=0\right\}$ (14)
$\mathcal{M}_{v}\triangleq\left\{(x,z,v)\in\mathbb{R}^{3}:v=0\right\}$ (15)
Considering the proof of Lemma 2.4, $\dot{\mathcal{V}}(0)=0$ if and only if $h_{2}(z)f_{1}(z)g_{1}(v)=0$, $h_{2}(z)f_{2}(z)g_{2}(v)=0$, and $vc(x,z,v)=0$. The following lemma can be inferred immediately :
Lemma 3.2.
$\mathcal{E}\subseteq\mathcal{M}_{v}\subseteq\dot{\mathcal{V}}^{-1}(0)\subseteq\mathcal{M}_{z}\cup\mathcal{M}_{v}$
Lemma 3.3.
With reference to the system given by Eq. (8), the largest invariant set that is contained in $\dot{\mathcal{V}}^{-1}(0)$ is $\mathcal{E}$.
Démonstration.
Due to Lemma 3.2, it suffices to show that the largest invariant set that is contained in $\mathcal{M}_{z}\cup\mathcal{M}_{v}$ is $\mathcal{E}$. Suppose that $X=(x,z,v):\mathbb{R}_{\geq 0}\longrightarrow\mathcal{M}_{z}\cup\mathcal{M}_{v}$ is a solution of the IVP given by Eq. (8).
Suppose that there exists $t_{s}\in\mathbb{R}_{\geq 0}$ such that
$X(t_{s})=(x(t_{s}),z(t_{s}),v(t_{s}))\in\mathcal{M}_{z}\setminus(\mathcal{M}_{z}\cap\mathcal{M}_{v})$
Then, $z(t_{s})=0$ and $v(t_{s})\neq 0$. By continuity of $X$, obtain $t_{e}\in\mathbb{R}_{>t_{s}}$ such that $X(t)\in\mathcal{M}_{z}\setminus(\mathcal{M}_{z}\cap\mathcal{M}_{v})$ for all $t\in[t_{s},t_{e})$. In this case, $z(t)=\dot{z}(t)=0$ for all $t\in[t_{s},t_{e})$. Therefore, $v(t)=0$ for all $t\in[t_{s},t_{e})$. However, in this case, $X(t)\in\mathcal{M}_{z}\cap\mathcal{M}_{v}$ for all $t\in[t_{s},t_{e})$, which results in a contradiction. Thus, $X(t)\subseteq\mathcal{M}_{v}$ for all $t\in\mathbb{R}_{\geq 0}$. Therefore, $v(t)=0$ for all $t\in\mathbb{R}_{\geq 0}$. Thus, $\dot{v}(t)=0$ for all $t\in\mathbb{R}_{\geq 0}$. Therefore, $h_{1}(x(t))+h_{2}(z(t))=0$ for all $t\in\mathbb{R}_{\geq 0}$. Thus, $X(t)\in\mathcal{E}\subseteq\mathcal{M}_{z}\cup\mathcal{M}_{v}$ for all $t\in\mathbb{R}_{\geq 0}$ or $X:\mathbb{R}_{\geq 0}\longrightarrow\mathcal{E}$. Since $\mathcal{E}$ is an invariant set, it is also the largest invariant set contained in $\mathcal{M}_{z}\cup\mathcal{M}_{v}$. ∎
Define the set $\mathcal{S}_{X}$ as
$\mathcal{S}_{X}\triangleq\{Y\in\mathbb{R}^{3}:\mathcal{V}(Y)\leq\mathcal{V}(X)\}$ (16)
Proposition 3.4.
Suppose that $X:\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}^{3}$ is a solution of the IVP given by Eq. (8) with $X_{0}\triangleq X(0)$. Then, $\mathcal{O}_{X_{0}}^{+\infty}$ is a nonempty, compact, connected, invariant set such that $\lim_{t\rightarrow+\infty}X(t)=\mathcal{O}_{X_{0}}^{+\infty}$ and $\mathcal{O}_{X_{0}}^{+\infty}\subseteq\mathcal{S}_{X_{0}}\cap\mathcal{E}$.
Démonstration.
Note that $\mathcal{O}_{X_{0}}^{+}$ is bounded by Proposition 2.5. Thus, $\mathcal{O}_{X_{0}}^{+\infty}$ is a nonempty, compact, connected, invariant set, and $\lim_{t\rightarrow+\infty}X(t)=\mathcal{O}_{X_{0}}^{+\infty}$ (e.g, see Proposition 5.1 in Ref. [bhat_nontangency-based_2003] or Theorem 2.41 in Ref. [haddad_nonlinear_2011]). Note that, by Lemma 3.3, the largest invariant set contained in $\dot{\mathcal{V}}^{-1}(0)$ is $\mathcal{E}$. Suppose that $\mathcal{P}$ is the largest invariant set contained in $\mathcal{S}_{X_{0}}$. Then, $\mathcal{O}_{X_{0}}^{+\infty}\subseteq\mathcal{P}\cap\mathcal{E}\subseteq\mathcal{S}_{X_{0}}\cap\mathcal{E}$ (e.g., see Proposition 5.3 in Ref. [bhat_nontangency-based_2003]). ∎
Define the set $\mathcal{E}_{xz}$ as
$\mathcal{E}_{xz}\triangleq\{(x,z)\in\mathbb{R}^{2}:h_{1}(x)=-h_{2}(z)\}$ (17)
The following corollary is an immediate consequence of Proposition 3.4 :
Corollary 3.5.
Suppose that $x,z,v:\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}$ form a solution of the IVP given by Eq. (8). Then, $\lim_{t\rightarrow+\infty}(x(t),z(t))=\mathcal{E}_{xz}$ and $\lim_{t\rightarrow+\infty}v(t)=0$.
Lemma 3.6.
Suppose that $H:\mathbb{R}\longrightarrow\mathbb{R}$ is a strictly decreasing continuous function such that $H(0)=0$. Define $J:\mathbb{R}\longrightarrow\mathbb{R}_{\geq 0}$ as $J(y)\triangleq d((0,y),\mathcal{G}(H))$ for all $y\in\mathbb{R}$. Then, $J$ is a continuous function such that $J(0)=0$ and $J(y)>0$ for all $y\in\mathbb{R}\setminus\{0\}$. Moreover, for every $y\in\mathbb{R}$, $J(y)=d((0,y),(a,b))$ for some $(a,b)\in\mathcal{G}(H)$ such that $a\leq 0$ and $0\leq b$ if $0\leq y$, and $0\leq a$ and $b\leq 0$ if $y\leq 0$. Furthermore, $J\restriction\mathbb{R}_{\geq 0}$ is nondecreasing and $J\restriction\mathbb{R}_{\leq 0}$ is nonincreasing.
Démonstration.
Define the function $G:\mathbb{R}\times\mathbb{R}\longrightarrow\mathbb{R}_{\geq 0}$ as
$G(y,x)\triangleq d((0,y),(x,H(x)))=\sqrt{x^{2}+(y-H(x))^{2}}$
and note that $(y)=\inf_{x\in\mathbb{R}}G(y,x). JJ$ is continuous due to continuity of $d(\cdot,S)$ for any non-empty set $S$. Since $H(0)=0$, $(0,0)\in\mathcal{G}(H)$, and, therefore, $J(0)=d((0,0),\mathcal{G}(H))=0$.
Suppose to the contrary that $J(y)=0$ for some $y\neq 0$. Without the loss of generality assume that $y>0$. Define
$\varepsilon\triangleq\min\left(-H^{-1}\left(\frac{y}{2}\right),\frac{y}{2}\right)$
[…]
For every $b\in\mathbb{R}_{>0}$, define $f_{b}^{+}:\mathbb{R}_{>0}\longrightarrow\mathbb{R}_{\geq 0}$ as
$f_{b}^{+}(y)\triangleq\min_{x\in[H^{-1}(b),0]}G(y,x)$
for every $y\in\mathbb{R}_{>0}$.
It will now be shown that $J(y)=f_{y}^{+}(y)$ for all $y\in\mathbb{R}_{>0}$. Fix $y\in\mathbb{R}_{>0}$. Note that $f_{y}^{+}(y)\leq y=G(y,0)$. Suppose that $x\in\mathbb{R}_{>0}$. Then, since $H(x)<0$,
[…]
Therefore,
$f_{y}^{+}(y)\leq y\leq\inf_{x\in(0,+\infty)}G(y,x)$
[…]
Thus,
[…]
Since $f_{y}^{+}(y)$, $\inf_{x\in(0,+\infty)}G(y,x)$, and $\inf_{x\in\left(-\infty,H^{-1}(y)\right)}G(y,x)$ are all bounded from below, $J(y)=f_{y}^{+}(y)$ by Lemma A.5. This also implies that
$J(y)=f_{b}^{+}(y)=\min_{x\in[H^{-1}(b),0]}(G\restriction(0,b)\times[H^{-1}(b),0])(y,x)$
for all $y\in(0,b)$ for all $b\in\mathbb{R}_{>0}$. Furthermore, this implies that, for every $y\in\mathbb{R}_{\geq 0}$, there exists $a\in[H^{-1}(y),0]$ such that $J(y)=d((0,y),(a,H(a)))$ (with $a\leq 0$ and $0\leq H(a)$).
Note that $J(y)<G(y,x)$ for all $y\in\mathbb{R}_{>0}$ and $x\in\mathbb{R}_{<H^{-1}(y)}$. To show this, fix $y\in\mathbb{R}_{>0}$ and $x\in\mathbb{R}_{<H^{-1}(y)}$ and, by the properties of $H$, obtain $z\in\mathbb{R}_{>0}$ such that $y<z$ and $x=H^{-1}(z)$. Then,
[…]
Fix $b\in\mathbb{R}_{>0}$. Noting that $G\restriction(0,b)\times[H^{-1}(b),0]$ is continuous and continuously differentiable with respect to the first argument, $(0,b)$ is open, $H^{-1}(b),0]$ is compact, by the Danskin’s Theorem ([danskin_theory_1966], see also Theorem 1.29 in Ref. [guler_foundations_2010]), the derivative of $J$ exists for every $y\in(0,b)$, and is given by
$\partial J(y)=\min_{x\in S(y)}\partial_{1}G(y,x)=\min_{x\in S(y)}\frac{y-H(x)}{\sqrt{x^{2}+(y-H(x))^{2}}}$
where
$S(y)\triangleq\{x\in[H^{-1}(b),0]:J(y)=G(y,x)\}$
Fix $y\in(0,b)$. Taking into account that $J(y)<G(y,x)$ for all $x\in\mathbb{R}_{<H^{-1}(y)}$, $0\leq y-H(x)$ for all $x\in S(y)$. Thus, $\partial J(y)\geq 0$ for all $y\in\mathbb{R}_{>0}$. Therefore, $J$ is nondecreasing on $\mathbb{R}_{>0}$. Since $J(0)=0$ and $J(y)>0$ for all $y\in\mathbb{R}_{>0}$, $J$ is nondecreasing on $\mathbb{R}_{\geq 0}$.
By similar arguments, it can be shown that for every $y\in\mathbb{R}_{\leq 0}$, there exists $a\in[0,H^{-1}(y)]$ such that $J(y)=d((0,y),(a,H(a)))$ (with $0\leq a$ and $H(a)\leq 0$), as well as that $J$ is nonincreasing on $\mathbb{R}_{\leq 0}$. ∎
By similar arguments it is possible to prove the following lemma :
Lemma 3.7.
Suppose that $H:\mathbb{R}\longrightarrow\mathbb{R}$ is a strictly decreasing continuous function such that $H(0)=0$. Define $J:\mathbb{R}\longrightarrow\mathbb{R}_{\geq 0}$ as $J(x)\triangleq d((x,0),\mathcal{G}(H))$ for all $x\in\mathbb{R}$. Then, $J$ is a continuous function such that $J(0)=0$ and $J(x)>0$ for all $x\in\mathbb{R}\setminus\{0\}$. Moreover, for every $x\in\mathbb{R}$, $J(x)=d((x,0),(a,b))$ for some $(a,b)\in\mathcal{G}(H)$ such that $a\leq 0$ and $0\leq b$ if $x\leq 0$, and $0\leq a$ and $b\leq 0$ if $0\leq x$. Furthermore, $J\restriction\mathbb{R}_{\geq 0}$ is nondecreasing and $J\restriction\mathbb{R}_{\leq 0}$ is nonincreasing.
Define the sets
$\mathcal{M}_{x}^{+}\triangleq\{(x,z,v)\in\mathbb{R}^{3}:z=0\wedge 0\leq x\}$ (18)
$\mathcal{M}_{x}^{-}\triangleq\{(x,z,v)\in\mathbb{R}^{3}:z=0\wedge x\leq 0\}$ (19)
$\mathcal{M}_{z}^{+}\triangleq\{(x,z,v)\in\mathbb{R}^{3}:x=0\wedge 0\leq z\}$ (20)
$\mathcal{M}_{z}^{-}\triangleq\{(x,z,v)\in\mathbb{R}^{3}:x=0\wedge z\leq 0\}$ (21)
$\mathcal{M}_{xz}\triangleq\mathcal{M}_{x}^{+}\cup\mathcal{M}_{x}^{-}\cup\mathcal{M}_{z}^{+}\cup\mathcal{M}_{z}^{-}$ (22)
Lemma 3.8.
Suppose that $X=(x,z,v):\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}^{3}$ is a solution of the IVP given by Eq. (8). Suppose that $X$ is in $\mathcal{M}_{z}^{+}$ frequently, that is, for all $T\in\mathbb{R}_{>0}$ there exists $t\in\mathbb{R}_{>T}$ such that $X(t)\in\mathcal{M}_{z}^{+}$. Then, $\lim_{t\rightarrow+\infty}X(t)=0$.
Démonstration.
Fix $\varepsilon\in\mathbb{R}_{>0}$. By Proposition 3.1, obtain $\delta\in\mathbb{R}_{>0}$ such that if $X(T)\in\mathbb{B}(0,\delta)$ for some $T\in\mathbb{R}_{\geq 0}$, then $X(t)\in\mathbb{B}(0,\varepsilon)$ for all $t>T$. Define $H:\mathbb{R}\longrightarrow\mathbb{R}$ as $H(x)\triangleq h_{2}^{-1}(-h_{1}(x))$ for all $x\in\mathbb{R}$. Note that $H$ is a strictly decreasing continuous function such that $H(0)=0$. Note also that $\mathcal{E}_{xz}=\mathcal{G}(H)$. Define $J:\mathbb{R}\longrightarrow\mathbb{R}_{\geq 0}$ as $J(z)\triangleq d((0,z),\mathcal{E}_{xz})$.
[…]
$d((x(t),z(t)),\mathcal{E}_{xz})=d((0,z(t)),\mathcal{E}_{xz})=J(z(t))<J\left(\delta/\sqrt{2}\right)$
Thus, $z(t)\leq\delta/\sqrt{2}$ by Lemma 3.6. Therefore,
$\lVert X(t)\rVert_{2}=\sqrt{x(t)^{2}+z(t)^{2}+v(t)^{2}}<\sqrt{\frac{\delta^{2}}{2}+\frac{\delta^{2}}{2}}=\delta$
Thus, $X(t)\in\mathbb{B}(0,\delta)$. Therefore, $X(s)\in\mathbb{B}(0,\varepsilon)$ for all $s\in\mathbb{R}_{>t}$. By generalization, for every $\varepsilon\in\mathbb{R}_{>0}$, there is $T\in\mathbb{R}_{\geq 0}$ such that $X(t)\in\mathbb{B}(0,\varepsilon)$ for all $t>T$. Therefore, $\lim_{t\rightarrow+\infty}X(t)=0$. ∎
The argument that was used in the proof of Lemma 3.8 can be applied to other half-planes :
Lemma 3.9.
Suppose that $X:\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}^{3}$ is a solution of the IVP given by Eq. (8). Suppose that $i\in\{x,z\}$ and $j\in\{-,+\}$. Suppose also that $X$ is in $\mathcal{M}_{i}^{j}$ frequently. Then, $\lim_{t\rightarrow+\infty}X(t)=0$.
Define
$\mathcal{N}_{ne}\triangleq\{(x,z,v)\in\mathbb{R}^{3}:x>0\wedge z>0\}$ (23)
$\mathcal{N}_{nw}\triangleq\{(x,z,v)\in\mathbb{R}^{3}:x<0\wedge z>0\}$ (24)
$\mathcal{N}_{sw}\triangleq\{(x,z,v)\in\mathbb{R}^{3}:x<0\wedge z<0\}$ (25)
$\mathcal{N}_{se}\triangleq\{(x,z,v)\in\mathbb{R}^{3}:x>0\wedge z<0\}$ (26)
$\mathcal{N}_{xz}\triangleq\mathcal{N}_{ne}\cup\mathcal{N}_{nw}\cup\mathcal{N}_{sw}\cup\mathcal{N}_{se}$ (27)
Lemma 3.10.
Suppose that $X:\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}^{3}$ is a solution of the IVP given by Eq. (8). Suppose also that $X$ is not in $\mathcal{M}_{xz}$ eventually, that is, there exists $T\in\mathbb{R}_{\geq 0}$ such that $X(t)\not\in\mathcal{M}_{xz}$ for all $t\in\mathbb{R}_{>T}$. Then, there exists $i\in\{ne,nw,sw,se\}$ such that $X$ is in $\mathcal{N}_{i}$ eventually.
Démonstration.
Obtain $T\in\mathbb{R}_{>0}$ such that $X(t)\not\in\mathcal{M}_{xz}$ for all $t\in\mathbb{R}_{>T}$. Therefore, $X(t)\in\mathbb{R}^{3}\setminus\mathcal{M}_{xz}=\mathcal{N}_{xz}$ for all $t\in\mathbb{R}_{>T}$. Note that $X(\mathbb{R}_{>T})$ is a connected set. Note also that $\mathcal{N}_{ne}$, $\mathcal{N}_{nw}$, $\mathcal{N}_{sw}$, and $\mathcal{N}_{se}$ are pairwise separated. Thus, by Lemma A.3, there is $i\in\{ne,nw,sw,se\}$ such that $X(t)\in\mathcal{N}_{i}$ for all $t\in\mathbb{R}_{>T}$. ∎
Lemma 3.11.
Suppose that $X:\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}^{3}$ is a solution of the IVP given by Eq. (8). Suppose also that $X$ is in $\mathcal{N}_{ne}$ eventually or $X$ is in $\mathcal{N}_{sw}$ eventually. Then, $\lim_{t\rightarrow+\infty}X(t)=0$.
Démonstration.
Fix $\varepsilon\in\mathbb{R}_{>0}$. By Proposition 3.1, obtain $\delta\in\mathbb{R}_{>0}$ such that if $X(T)\in\mathbb{B}_{\infty}(0,\delta)$ for some $T\in\mathbb{R}_{\geq 0}$, then $X(t)\in\mathbb{B}(0,\varepsilon)$ for all $t>T$. Suppose that $X$ is in $\mathcal{N}_{ne}$ eventually. Obtain $T_{1}\in\mathbb{R}_{\geq 0}$ such that $X(t)\in\mathcal{N}_{ne}$ for all $t\in\mathbb{R}_{>T_{1}}$. Define $H:\mathbb{R}\longrightarrow\mathbb{R}$ as $H(x)\triangleq h_{2}^{-1}(-h_{1}(x))$ for all $x\in\mathbb{R}$. Note that $H$ is a strictly decreasing continuous function such that $H(0)=0$. Note also that $\mathcal{E}_{xz}=\mathcal{G}(H)$. Define $J_{1}:\mathbb{R}\longrightarrow\mathbb{R}_{\geq 0}$ as $J_{1}(z)\triangleq d((0,z),\mathcal{E}_{xz})$ and $J_{2}:\mathbb{R}\longrightarrow\mathbb{R}_{\geq 0}$ as $J_{2}(x)\triangleq d((x,0),\mathcal{E}_{xz})$.
[…]
$\mathcal{E}_{xz}^{+}\triangleq\{(x,H(x)):x\in\mathbb{R}_{\leq 0}\}$
$\mathcal{E}_{xz}^{-}\triangleq\{(x,H(x)):x\in\mathbb{R}_{\geq 0}\}$
and note that $\mathcal{E}_{xz}=\mathcal{E}_{xz}^{+}\cup\mathcal{E}_{xz}^{-}$. Then, by Lemma A.5, either
$d((x(t),z(t)),\mathcal{E}_{xz})=d((x(t),z(t)),\mathcal{E}_{xz}^{+})$
or
$d((x(t),z(t)),\mathcal{E}_{xz})=d((x(t),z(t)),\mathcal{E}_{xz}^{-})$
Suppose that $\delta\leq z(t)$. Suppose also that
$d((x(t),z(t)),\mathcal{E}_{xz})=d((x(t),z(t)),\mathcal{E}_{xz}^{-})$
Then,
$d((x(t),z(t)),\mathcal{E}_{xz})$ $=\inf_{(a,b)\in\mathcal{E}_{xz}^{-}}\sqrt{(x(t)-a)^{2}+(z(t)-b)^{2}}$
[…]
$>d((x(t),z(t)),\mathcal{E}_{xz})$
Therefore, by contradiction,
$d((x(t),z(t)),\mathcal{E}_{xz})=d((x(t),z(t)),\mathcal{E}_{xz}^{+})$
Noting that
$d((x(t),z(t)),\mathcal{E}_{xz}^{+})\geq d((0,z(t)),\mathcal{E}_{xz}^{+})=d((0,z(t)),\mathcal{E}_{xz})$
[…]
The proof follows by a similar argument if $X$ is in $\mathcal{N}_{sw}$ eventually. ∎
Define $\mathcal{W}:\mathbb{R}^{3}\longrightarrow\mathbb{R}$ as
$\mathcal{W}(x,z,v)\triangleq(z-x)^{2}$ (28)
for all $x,z,v\in\mathbb{R}$. Note that
$\dot{\mathcal{W}}(x,z,v)=2(z-x)(f_{1}(z)g_{1}(v)+f_{2}(z)g_{2}(v))$ (29)
for all $x,z,v\in\mathbb{R}$. Define
$\mathcal{A}(L)\triangleq\{(x,z,v)\in\mathbb{R}^{3}:z=x+L\wedge v=0\}$ (30)
for all $L\in\mathbb{R}$.
Lemma 3.12.
Suppose that $X:\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}^{3}$ is a solution of the IVP given by Eq. (8). If $X$ is in $\mathcal{N}_{nw}$ eventually, then there exists $L\in\mathbb{R}_{\geq 0}$ such that $\lim_{t\rightarrow+\infty}X(t)=\mathcal{A}(L)$. If $X$ is in $\mathcal{N}_{se}$ eventually, then there exists $L\in\mathbb{R}_{\leq 0}$ such that $\lim_{t\rightarrow+\infty}X(t)=\mathcal{A}(L)$.
Démonstration.
[…]
$<\sqrt{\frac{1}{2}\varepsilon^{2}+\frac{1}{2}\varepsilon^{2}}=\varepsilon$
By generalization, for all $\varepsilon\in\mathbb{R}_{>0}$, there exists $T\in\mathbb{R}_{>0}$ such that $d((x(t),z(t),v(t)),\mathcal{A}(L))<\varepsilon$. Thus, $\lim_{t\rightarrow+\infty}X(t)=\mathcal{A}(L)$.
The proof is similar if $X$ is in $\mathcal{N}_{se}$ eventually. ∎
Lemma 3.13.
Suppose that $X:\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}^{3}$ is a solution of the IVP given by Eq. (8) with $X_{0}\triangleq X(0)$. Suppose also that $X$ is in $\mathcal{N}_{nw}$ eventually or $X$ is in $\mathcal{N}_{se}$ eventually. Then, there exists $X^{*}\in\mathcal{E}\cap\mathcal{S}_{X_{0}}$ such that $\lim_{t\rightarrow+\infty}X(t)=X^{*}$.
Démonstration.
Define $H:\mathbb{R}\longrightarrow\mathbb{R}$ as $H(x)\triangleq h_{2}^{-1}(-h_{1}(x))$ for all $x\in\mathbb{R}$. Note that $H$ is a strictly decreasing continuous function such that $H(0)=0$.
Suppose that $X$ is in $\mathcal{N}_{nw}$ eventually. Obtain $T_{1}\in\mathbb{R}_{>0}$ such that $X(t)\in\mathcal{N}_{nw}$ for all $t\in\mathbb{R}_{>T_{1}}$. By Lemma 3.12, obtain $L\in\mathbb{R}_{\geq 0}$ such that $\lim_{t\rightarrow+\infty}X(t)=\mathcal{A}(L)$.
By Lemma A.4, obtain the unique $a\in\mathbb{R}_{\leq 0}$ such that
$X^{*}\triangleq(a,H(a),0)=(a,a+L,0)$
Then, $\mathcal{E}\cap\mathcal{A}(L)=\{X^{*}\}$.
Fix $p\in\mathcal{O}_{X_{0}}^{+\infty}$. Fix $\varepsilon\in\mathbb{R}_{>0}$. Note that $\mathcal{O}_{X_{0}}^{+\infty}\subseteq\mathcal{E}\cap\text{cl}\>\mathcal{N}_{nw}$ with $\mathcal{O}_{X_{0}}^{+\infty}$ being a compact set by Proposition 3.4. Obtain $T_{2}\in\mathbb{R}_{>T_{1}}$ such that $d(X(t),\mathcal{A}(L))<\varepsilon/2$ for all $t\in\mathbb{R}_{>T_{2}}$. Obtain $t\in\mathbb{R}_{>T_{2}}$ such that $d(p,X(t))<\varepsilon/2$. Then, by the triangle inequality,
$d(p,\mathcal{A}(L))\leq d(p,X(t))+d(X(t),\mathcal{A}(L))<\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon$
Thus, by generalization, $d(p,\mathcal{A}(L))<\varepsilon$ for all $\varepsilon\in\mathbb{R}_{\geq 0}$. Therefore, by analysis, $d(p,\mathcal{A}(L))=0$. Since $\mathcal{A}(L)$ is a closed subset of $\mathbb{R}^{3}$, $p\in\mathcal{A}(L)$. Thus, by generalization, $\mathcal{O}_{X_{0}}^{+\infty}\subseteq\mathcal{A}(L)$. Thence,
$\mathcal{O}_{X_{0}}^{+\infty}=\mathcal{O}_{X_{0}}^{+\infty}\cap\mathcal{A}(L)\subseteq\mathcal{E}\cap\mathcal{A}(L)=\{X^{*}\}$
Therefore, $\lim_{t\rightarrow+\infty}X(t)=X^{*}$.
The proof is similar if $X$ is in $\mathcal{N}_{se}$ eventually. ∎
Theorem 3.14.
Suppose that $X:\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}^{3}$ is a solution of the IVP given by Eq. (8) with $X_{0}\triangleq X(0)$. Then, there exists $X^{*}\in\mathcal{E}\cap\mathcal{S}_{X_{0}}$ such that $\lim_{t\rightarrow+\infty}X(t)=X^{*}$.
Démonstration.
Suppose that $X$ is in $\mathcal{M}_{xz}$ frequently. Then, $\lim_{t\rightarrow+\infty}X(t)=0\in\mathcal{E}\cap\mathcal{S}_{X_{0}}$ by Lemma 3.9. Suppose that $X$ is not in $\mathcal{M}_{xz}$ eventually. Then, by Lemma 3.10, there exists $i\in\{ne,nw,sw,se\}$ such that $X$ is in $\mathcal{N}_{i}$ eventually. Suppose that $X$ is in $\mathcal{N}_{ne}$ eventually or $X$ is in $\mathcal{N}_{sw}$ eventually. Then, by Lemma 3.11, $\lim_{t\rightarrow+\infty}X(t)=0\in\mathcal{E}\cap\mathcal{S}_{X_{0}}$. Suppose that $X$ is in $\mathcal{N}_{nw}$ eventually or $X$ is in $\mathcal{N}_{se}$ eventually. Then, by Lemma 3.13, there exists $X^{*}\in\mathcal{E}\cap\mathcal{S}_{X_{0}}$ such that $\lim_{t\rightarrow+\infty}X(t)=X^{*}$. ∎
4 The Bouc-Wen Oscillator
In the context of plasticity, arguably, the most well-known specialization of the Duhem-models is the Bouc-Wen model [bouc_forced_1967, bouc_modemathematique_1971, wen_method_1976]. The Bouc-Wen model is a general parameterizable rate-independent differential model of hysteresis. It can be described by the following system of differential equations with an input and an output [ikhouane_systems_2007]
[…]
where $x,z\in\mathbb{R}$ are state variables, $v\in\mathbb{R}$ is an input variable, $F\in\mathbb{R}$ is an output variable, $A,\beta,\gamma\in\mathbb{R}$, $\alpha\in(0,1)\subseteq\mathbb{R}$, $k\in\mathbb{R}_{>0}$, $D\in\mathbb{R}_{>0}$, $n\in\mathbb{R}_{>1}$, $\beta\neq-\gamma$, $x_{0},z_{0}\in\mathbb{R}$ are parameters. Recent surveys in Refs. [ikhouane_systems_2007, ismail_hysteresis_2009, heredia-perez_state---art_2025] provide further general information about the model.
According to Ref. [ikhouane_dynamic_2007], the Bouc-Wen model is of class I if and only if $A\in\mathbb{R}_{>0}$, $\gamma\in(-\beta,\beta]$. The authors of Ref. [ikhouane_dynamic_2007] describe other four distinct classes of the Bouc-Wen model that depend on the model parameters. However, these classes are seldom relevant for the applications related to mechanical oscillations : “class I is the only one that is BIBO stable, is compatible with the free motion of the real systems described by the Bouc–Wen model, is passive and is compatible with the laws of thermodynamics” [ikhouane_dynamic_2007].
In Ref. [ikhouane_dynamic_2007], the authors study a model of a point mass attached to a rigid surface via a viscous damper and a Bouc-Wen elastoplastic element. The dynamics of the system are given by the following IVP :
[…]
Here, $x,z,v\in\mathbb{R}$ are state variables, $m\in\mathbb{R}_{>0}$, $b\in\mathbb{R}_{\geq 0}$, and $v_{0}\in\mathbb{R}$ are additional model parameters. In Ref. [ikhouane_dynamic_2007], it is shown that if the Bouc-Wen model belongs to class I or class II, then there exists $\bar{b}\in\mathbb{R}_{\geq 0}$ such that $\bar{b}\leq b$ implies that $x_{\infty},z_{\infty}\in\mathbb{R}$ such that $\lim_{t\rightarrow+\infty}x(t)=x_{\infty}$, $\lim_{t\rightarrow+\infty}z(t)=z_{\infty}$, $\lim_{t\rightarrow+\infty}v(t)=0$, and $\alpha x_{\infty}+(1-\alpha)Dz_{\infty}=0$.
The Bouc-Wen oscillator is a special case of the Duhem oscillator that was considered in this study. To show this, introduce a new state variable $Z\triangleq z/(AD^{-1})$. Then, after rescaling and renaming (i.e., replacing the symbol $Z$ with $z$), Eq. (32) becomes
[…]
Define $f_{1},f_{2},g_{1},g_{2},h_{1},h_{2}:\mathbb{R}\longrightarrow\mathbb{R}$ and $c:\mathbb{R}^{3}\longrightarrow\mathbb{R}$ as
[…]
$h_{1}(x)\triangleq\alpha\frac{k}{m}x$ (38)
$h_{2}(z)\triangleq(1-\alpha)A\frac{k}{m}z$ (39)
$c(x,z,v)\triangleq\frac{b}{m}v$ (40)
Then, Eq. (33) can be written as
$\begin{cases}\dot{x}=v\\ \dot{z}=v+f_{1}(z)g_{1}(v)+f_{2}(z)g_{2}(v)\\ \dot{v}=-h_{1}(x)-h_{2}(z)-c(x,z,v)\end{cases}$ (41)
which is equivalent to Eq. (8). It can be verified that the functions $f_{1}$, $f_{2}$, $g_{1}$, $g_{2}$, $h_{1}$, $h_{2}$, $c$ satisfy the conditions necessary for the application of Theorem 3.14 provided that $A\in\mathbb{R}_{>0}$, $\gamma\in(-\beta,\beta]$, and $b\in\mathbb{R}_{\geq 0}$ (the Bouc-Wen model of class I). In this case, by Theorem 3.14, there exist $x_{\infty},z_{\infty}\in\mathbb{R}$ such that $\lim_{t\rightarrow+\infty}x(t)=x_{\infty}$, $\lim_{t\rightarrow+\infty}z(t)=z_{\infty}$, $\lim_{t\rightarrow+\infty}v(t)=0$, and $\alpha x_{\infty}+(1-\alpha)Az_{\infty}=0$.
5 Conclusions and Future Work
The article described certain analytical properties of an unforced mechanical oscillator with a Duhem-type viscoelastoplastic hysteretic element. These properties include convergence of each solution to an equilibrium point, generalizing a result that was previously presented in Ref. [ikhouane_dynamic_2007].
Future work may include generalization of the results presented in this article to other (more abstract) Duhem-type models, an investigation of the stability of individual equilibrium points, and estimation of the rate of convergence of the solutions.
Acknowledgment
The authors would like to acknowledge their families, colleagues, and friends. Special thanks go to the members of staff of Auburn University Libraries for their assistance in finding rare and out-of-print research articles and research monographs. The authors would also like to acknowledge the professional online communities, instructional websites, and various online service providers, especially https://www.adobe.com/acrobat/online/pdf-to-word.html, https://archive.org/, https://automeris.io, https://capitalizemytitle.com, https://www.matweb.com, https://www.overleaf.com, https://pgfplots.net, https://proofwiki.org/, https://www.reddit.com, https://scholar.google.com, https://stackexchange.com, https://stringtranslate.com, https://www.wikipedia.org. We also note that the results of some of the calculations that are presented in this article were performed with the assistance of the software Wolfram Mathematica [wolfram_research_inc_mathematica_2023]. Other software that was used to produce this article included Adobe Acrobat Reader, Adobe Digital Editions, DiffMerge, Git, GitLab, Google Chrome, Google Gemini (Google Gemini was used as an assistant ; no significant parts of the article were written by AI), Grammarly (the use of Grammarly was restricted to the identification and correction of spelling, grammar, and punctuation errors), Jupyter Notebook, LibreOffice, macOS Monterey, Mamba, Microsoft Outlook, Preview, Safari, TeX Live/MacTeX, Texmaker, and Zotero.
Funding Data
The present work did not receive any specific funding. However, the researchers receive financial support from Auburn University for their overall research activity.
Annexe A Notation, Conventions, Foundations
Essentially all of the definitions and results that are employed in this article are standard in the fields of set theory, general topology, analysis, ordinary differential equations, and nonlinear systems/control. They can be found in a number of textbooks and monographs on these subjects (e.g., see Ref. [takeuti_introduction_1982], Refs. [kelley_general_1955, morris_topology_2020, baldwin_math_2024], Refs. [bloch_real_2010, shurman_calculus_2016, ziemer_modern_2017], Refs. [chicone_ordinary_1999, schaeffer_ordinary_2016], Refs. [lasalle_extensions_1960, yoshizawa_stability_1966, yoshizawa_stability_1975, sontag_mathematical_1998, sastry_nonlinear_1999, khalil_nonlinear_2002, haddad_nonlinear_2011], respectively).
Definition A.1.
$\in$ denotes the set membership relation, $\subseteq$ denotes the subset relation, $\subset$ denotes the proper subset relation, $\cup$ denotes the binary set union operation, $\cap$ denotes the binary set intersection operation, $\setminus$ denotes the binary set difference operation, $\mathcal{P}$ denotes the power set operation, $\emptyset$ denotes the empty set, $(a_{1},\ldots,a_{n})$ denotes an $n$-tuple, $\{a_{1},\ldots,a_{n}\}$ denotes an unordered collection of elements. [^4] It should be noted that some of the syntactic constructions may carry different semantics depending on the context. For example, $(a,b)$ may be used as a pair or as an interval. It is hoped that the context of the discussion will always make the meaning of a given syntactic construction apparent.
Definition A.2.
By convention, a topological space cannot be empty. Suppose $X\neq\emptyset$ and $\tau\subseteq\mathcal{P}X$ is a topology on $X$. $\mathsf{cl}A$ denotes the closure of $A\subseteq X$ ; if $Y\subseteq X$ and $Y\neq\emptyset$, then $\tau|Y$ will denote the subspace topology of $\tau$ on $Y$ ; the sets $A\subseteq X$ and $B\subseteq X$ are separated if and only if $\mathsf{cl}A\cap B=A\cap\mathsf{cl}B=\emptyset$ ; a set $C\subseteq X$ is clopen if and only if it is open and closed ; $A\subseteq X$ is connected if and only if it is not a union of two nonempty separated sets ; $(X,\tau)$ is a connected topological space if and only if $X$ is a connected set.
It should be noted that different definitions of a connected set and a connected topological space are employed in some of the cited literature. The following technical lemmas establish a connection between the two commonly used definitions (these results are not used directly, and the proofs were deemed to be sufficiently simple to be omitted) :
Lemma A.1.
Suppose $(X,\tau)$ is a topological space. Then, $(X,\tau)$ is connected if and only if the only clopen sets in $(X,\tau)$ are $\emptyset$ and $X$.
Lemma A.2.
Suppose $(X,\tau)$ is a topological space and $Y\subseteq X$. Then, $Y$ is a connected set in $(X,\tau)$ if and only if either $Y=\emptyset$ or $(Y,\tau|Y)$ is a connected topological space.
The proof of the following lemma was deemed to be sufficiently simple to be omitted :
Lemma A.3.
Suppose $(X,\tau)$ is a topological space. Suppose that $A\subseteq X$ and $B\subseteq X$ are separated, $C\subseteq A\cup B$ is connected. Then, $C\subseteq A$ or $C\subseteq B$.
Definition A.3.
$\mathbb{Z}$ is the set of all integers ; $\mathbb{R}$ is the set of all real numbers ; an interval of real numbers $I\subseteq\mathbb{R}$ is non-degenerate if it has a non-empty interior ; $\mathbb{K}_{>a}\triangleq(a,+\infty)\cap\mathbb{K}$, $\mathbb{K}_{<a}\triangleq(-\infty,a)\cap\mathbb{K}$, $\mathbb{K}_{\geq a}\triangleq[a,+\infty)\cap\mathbb{K}$, and $\mathbb{K}_{\leq a}\triangleq(-\infty,a]\cap\mathbb{K}$ for any $a\in\mathbb{R}$ with $\mathbb{K}\subseteq\mathbb{R}$ ; $\mathbb{R}^{n}$ with $n\in\mathbb{Z}_{\geq 1}$ is the set of $n$-tuples of real numbers (augmented with the structure of the Euclidean space) ; if $X=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}$ with $n\in\mathbb{Z}_{\geq 1}$, then $X_{i}\triangleq x_{i}$ for all $i\in\{1,\ldots,n\}$ ; $f:X\longrightarrow Y$ denotes a function with the domain $X$ and the codomain $Y$ ; given $f:X\longrightarrow Y$, $f(A)$ denotes the image of $f$ under the set $A$ ; if $f:X\longrightarrow Y$, then $\mathcal{G}(f)\subseteq X\times Y$ is the graph of $f$ given by $\mathcal{G}(f)\triangleq\{(x,y)\in X\times Y:y=f(x)\}$ ; if $f:X\longrightarrow\mathbb{R}^{n}$ with $n\in\mathbb{Z}_{\geq 1}$, then $f_{i}:X\longrightarrow\mathbb{R}$ is given by $f_{i}(x)\triangleq(f(x))_{i}$ for all $x\in X$ and $i\in\{1,\ldots,n\}$ ; unless stated otherwise, the topology of a subset of $\mathbb{R}^{n}$ with $n\in\mathbb{Z}_{\geq 1}$ is always the subspace topology of the standard topology on $\mathbb{R}^{n}$ ; given $A\subseteq\mathbb{R}$, $\inf A\in\mathbb{R}\cup\{-\infty,+\infty\}$ denotes the infimum of $A$ and $\sup A\in\mathbb{R}\cup\{-\infty,+\infty\}$ denotes the supremum of $A$ ; given a sequence $\{x_{i}\in\mathbb{R}^{n}\}_{i\in\mathbb{Z}_{\geq 1}}$ with $n\in\mathbb{Z}_{\geq 1}$, $\lim_{i\rightarrow+\infty}x_{i}$ denotes the limit of $x$, provided that it exists ; $\langle\cdot,\cdot\rangle:\mathbb{R}^{n}\times\mathbb{R}^{n}\longrightarrow\mathbb{R}$ with $n\in\mathbb{Z}_{\geq 1}$ is the canonical inner product on $\mathbb{R}^{n}$ ; $\lVert\cdot\rVert_{p}:\mathbb{R}^{n}\longrightarrow\mathbb{R}_{\geq 0}$ with $n\in\mathbb{Z}_{\geq 1}$ and $p\in\mathbb{R}_{\geq 1}\cup\{+\infty\}$ is the $p$-norm on $\mathbb{R}^{n}$ ; given $n\in\mathbb{Z}_{\geq 1}$ and $p\in\mathbb{R}_{\geq 1}\cup\{+\infty\}$, $d_{p}:\mathbb{R}^{n}\times\mathbb{R}^{n}\longrightarrow\mathbb{R}_{\geq 0}$ given by $d_{p}(x,y)\triangleq\lVert x-y\rVert_{p}$ is the metric induced by the $p$-norm ; $d\triangleq d_{p}$ for all $p\in\mathbb{R}_{\geq 1}\cup\{+\infty\}$ ; assuming that $n\in\mathbb{Z}_{\geq 1}$, $p\in\mathbb{R}_{\geq 1}\cup\{+\infty\}$, $a\in\mathbb{R}^{n}$, and $r\in\mathbb{R}_{>0}$, $\mathbb{B}_{p}(a,r)\triangleq\{\ x\in\mathbb{R}^{n}:d_{p}(x,a)<r\}$ is an open $p$-ball in $\mathbb{R}^{n}$ centered at $a$ with the radius $r$ ; $f:\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n}$ with $n\in\mathbb{Z}_{\geq 1}$ is locally Lipschitz if and only if for every $x\in\mathbb{R}^{n}$ there exists an open set $U\subseteq\mathbb{R}^{n}$ such that $x\in U$ and there exists $L\in\mathbb{R}_{>0}$ such that $d(f(y),f(z))\leq Ld(y,z)$ for all $y,z\in U$ ; given a differentiable function $f:X\longrightarrow Y$ such that $X\subseteq\mathbb{R}$ and $Y\subseteq\mathbb{R}^{n}$ with $n\in\mathbb{Z}_{\geq 1}$, $df/dx$ and $\partial f$ may be used to denote the derivative of the function ; the overdot notation $\dot{x}\triangleq(dx/dt)$ may be used to represents the derivative of a differentiable function $x:\mathbb{R}\longrightarrow\mathbb{R}^{n}$ with $n\in\mathbb{Z}_{\geq 1}$ with respect to the time variable in the context of mechanics ; given a differentiable function $f:X\longrightarrow Y$ such that $X\subseteq\mathbb{R}^{n}$ and $Y\subseteq\mathbb{R}$ with $n\in\mathbb{Z}_{\geq 1}$, $\partial_{i}f$ denotes the $i$-th partial derivative of the function for $i\in\{1,\ldots,n\}$.
Lemma A.4.
Suppose that $f:\mathbb{R}\longrightarrow\mathbb{R}$ is a continuous strictly decreasing function such that $f(0)=0$. Suppose that $g:\mathbb{R}\longrightarrow\mathbb{R}$ is a continuous strictly increasing function such that $g(a)=0$ and $g(0)=b$ for some $a\in\mathbb{R}_{\leq 0}$ and $b\in\mathbb{R}_{\geq 0}$. Then, there is a unique $x\in\mathbb{R}$ such that $f(x)=g(x)$. Moreover, $x\in\mathbb{R}_{\leq 0}$.
Démonstration.
Define the continuous function $h:\mathbb{R}\longrightarrow\mathbb{R}$ as
$h(x)\triangleq f(x)-g(x)$
for all $x\in\mathbb{R}$. Note that $h$ is a strictly decreasing continuous function. Then, $x=h^{-1}(0)\in\mathbb{R}$ is the unique point such that $h(x)=0$ or, equivalently, $f(x)=g(x)$. It remains to show that $x\in\mathbb{R}_{\leq 0}$.
Suppose that $a=0$. Then, $g(0)=0$. Since $f(0)=0$ by assumptions of the lemma, $x=0\in\mathbb{R}_{\leq 0}$ is such that $f(x)=g(x)$. Suppose that $b=0$. Then, $g(0)=0$. Since $f(0)=0$ by assumptions of the lemma, $x=0\in\mathbb{R}_{\leq 0}$ is such that $f(x)=g(x)$.
Suppose that $a<0$ and $b>0$. Note that h(a)=f(a)g(a)=f(a)>0h(a)=f(a)-g(a)=f(a)>0. Note also that h(0)=f(0)g(0)=b<0h(0)=f(0)-g(0)=-b<0. By the Intermediate Value Theorem (e.g., see Theorem 3.5.2 in Ref. [bloch_real_2010]), obtain $x\in(a,0)\subseteq\mathbb{R}_{\leq 0}$ such that $h(x)=0$. ∎
Lemma A.5.
Suppose that $A,B\subseteq\mathbb{R}$ are bounded from below and $\inf A\leq\inf B$. Then, $\inf(A\cup B)=\inf A$.
Démonstration.
It is apparent that $\inf(A\cup B)\leq\inf A$. Suppose that
$c\triangleq\inf A\cup B<\inf A\triangleq a$
Obtain $b\in A\cup B$ such that $b\in[c,a)$. Then, $b<\inf A\leq\inf B$, which results in a contradiction. Thus, $\inf(A\cup B)=\inf A$. ∎
Definition A.4.
Consider the following system of ordinary differential equations
$\dot{x}=f(x)\\$ (42)
where $f:\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n}$ with $n\in\mathbb{Z}_{\geq 1}$ is a locally Lipschitz continuous state function. Equation (42) augmented with an initial condition $x(0)=x_{0}\in\mathbb{R}^{n}$ shall be referred to as an Initial Value Problem (IVP) associated with the system given by Eq. (42). A differentiable function $x:I\longrightarrow\mathbb{R}^{n}$ with $I\subseteq\mathbb{R}$ being a non-degenerate interval such that $0\in I$ is a solution of the IVP associated with the system given by Eq. (42) with the initial condition $x_{0}\in\mathbb{R}^{n}$ if $x(0)=x_{0}$ and $\dot{x}(t)=f(x(t))$ for all $t\in I$.
The following definitions can be found in Ref. [bhat_nontangency-based_2003] and Ref. [haddad_nonlinear_2011] :
Definition A.5.
For the remainder of this definition, suppose that the system given by Eq. (42) has a unique solution defined on $\mathbb{R}_{\geq 0}$ for every initial condition. Suppose that $x:\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}^{n}$ is a solution of an IVP associated with the system given by Eq. (42) with the initial condition $x(0)=z\in\mathbb{R}^{n}$. Then, $\mathcal{O}_{z}^{+}\triangleq\{x(t):t\in\mathbb{R}_{\geq 0}\}$ is the positive orbit of $z$. A set $U\subseteq\mathbb{R}^{n}$ is positively invariant with respect to the system given by Eq. (42) if and only if for every solution $x:\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}^{n}$ of the IVP with $x(0)=z\in U$, $x(t)\in U$ for all $t\in\mathbb{R}_{\geq 0}$. A set $U\subseteq\mathbb{R}^{n}$ is negatively invariant with respect to the system given by Eq. (42) if and only if for every $z\in U$ and $T\in\mathbb{R}_{\geq 0}$ there exists a solution $x:[0,T]\longrightarrow U$ of the IVP with $x(T)=z$. A set $U\subseteq\mathbb{R}^{n}$ is invariant with respect to the system given by Eq. (42) if and only if it is positively invariant and negatively invariant with respect to the system given by Eq. (42). Suppose again that $x:\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}^{n}$ is a solution of an IVP associated with the system given by Eq. (42) with the initial condition $x(0)=z\in\mathbb{R}^{n}$. Then, $p\in\mathbb{R}^{n}$ is a positive limit point of $z$ if and only if there exists a nondecreasing sequence $\{t_{n}\}_{n\in\mathbb{Z}_{\geq 1}}$ of positive real numbers such that $\lim_{n\rightarrow+\infty}t_{n}=+\infty$ and $\lim_{n\rightarrow+\infty}x(t_{n})=p$. Furthermore, $\mathcal{O}_{z}^{+\infty}\subseteq\mathbb{R}^{n}$ shall be used to denote the positive limit set of $z$, that is, the set of all positive limit points of $z$. $\lim_{t\rightarrow+\infty}x(t)=A\subseteq\mathbb{R}^{n}$ if and only if for every $\varepsilon\in\mathbb{R}_{>0}$ there exists $T\in\mathbb{R}_{>0}$ such that $\inf_{p\in A}d(x(t),p)<\varepsilon$ for all $t\in\mathbb{R}_{>T}$. A continuous and strictly increasing function $\alpha:\mathbb{R}_{\geq 0}\longrightarrow\mathbb{R}_{\geq 0}$ is of class $\mathcal{K}_{\infty}$ if and only if $\alpha(0)=0$ and $\lim_{x\rightarrow+\infty}\alpha(x)=+\infty$.
The following definition can be found in Ref. [yoshizawa_stability_1975] :
Definition A.6.
The solutions of the system given by Eq. (42) are said to be equibounded if and only if for all $\alpha\in\mathbb{R}_{>0}$ there exists $\beta\in\mathbb{R}_{>0}$ such that $\lVert x(t)\rVert_{2}<\beta$ for all $t\in[0,T)$ for every solution $x:[0,T)\longrightarrow\mathbb{R}^{n}$ with $T\in\mathbb{R}_{>0}\cup\{+\infty\}$ starting from the initial condition $x(0)=x_{0}\in\mathbb{R}^{n}$ such that $\lVert x_{0}\rVert_{2}\leq\alpha$.